SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (, 2007)

Attendees: Nicolas Arnaud, Debbie Bard, Rainer Bartoldus, Brandon Drummond, Al Eisner, Fang Fang, Kim Hojeong, Arik Kreisel, Rahmat Rahmat.

Operations

  • Quiet this week! Please look over the slides for the quarterly trigger status report tomorrow.

EMT

  • During the ROD last week Mark swapped in another spare TPB. While taking the boards in/out of the crate some chips got knocked off the back of the board, and off the good board that was in the crate. The broken boards will be taken to Ray or Lupe for repair, but there may not be any spare chips for replacement! The good spare is in the crate for now.

GLT

  • During the ROD, Fang took cosmics run 70930 with the new GLT libraries in. The data looked fine, compared to the run 70929 reference, and the power-cycle problem was fixed with these libraries (no need to reboot the ROM). More detailed study is needed of the GLT data, once verified a physics run can be taken with the new libraries and that validated by the DQG.

L3

  • David Brown in the tracking group showed the improvement gained using the L3 tracking algorithms with TrkFixup - 10% reduction in ghost and bad tracks, 1% increase in good tracks. More MC will be generated to give them more statistics to validate this further, but it's a big success! The process will be summerised at a Wednesday meeting in ~2 weeks.

Data Quality and Monitoring

  • Those peaks in the bhabha/mumu pairs lumi srtipcharts Rahmat showed last week were NOT due to the introduction of new bhabha constants. He ran over reprocessed data and still saw this effect. He will produce a list of runs that show the higher ratio - it may be a prescaling effect? There's no problems with other quantities so it's not a physics problem.
  • The L3TSlowMon plots are not regularly looked at at present, they could be incorporated into the logbook or at least better presented.

Config

  • New MC is in progress, but the web page needs updating and presentation improved. Not all mode numbers are given, and not modes have aug06 conditions.

Trigger rate, deadtime and lumi monitoring

  • Nicolas showed some plots here which could be included in the weekly backgrounds/ run coordinators report.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.